Recent Log Entries

1L24 HV Drop - Solenoid Low

Log Number: 4258583Logbook(s): Entry Makers: dchriBody: 

Attempted to reset HV with no luck
Took it out of FIL, tried to reset HV, no luck
Bypassed the problem cavity (cav 6), tried to reset HV, no luck
In the mist of Contacting RF on call and unbypassing cav 6, Solenoid came back on momentarily
After attempting to bring HV and Solenoid tripped again
Decided to bypass cavity 6 again, retry to bring back HV, no luck

QE Measurement

Log Number: 4258581Logbook(s): Body: Spot Location: (12000,18000), Charge: 54.92293699999999

Laser Wavelength Atten Setting Laser Power Current HV PS Current QE PV Updated? Update Time/Reason Not Updated (nm) (W) (uA)</T

FSD Masking Tool: Successful System Masking for: CryoModule

Log Number: 4258577Logbook(s): Tags: Body: System Masking for: CryoModule
User: deir
Areas: All

CryoModule Devices:
0L02
0L04
0L03
1L14
1L03
1L02
1L04
1L05
1L07
1L06
1L08
1L10
1L11
1L09
1L12
1L15
1L13
1L16
1L19
1L17
1L18
1L20
1L21
1L22
1L23
1L24
1L25
1L26
2L14
2L03
2L02
2L04
2L07
2L05
2L06
2L08
2L11
2L09
2L10
2L12
2L15
2L13
2L16
2L19
2L17
2L18
2L20
2L21
2L22
2L23
2L24
2L25
2L26


xB Elevated Upstream Halo Rates

Log Number: 4258551Logbook(s): Entry Makers: deirBody: 

This is similar to the issue that arose this past Tuesday. Day shift did some steering throughout the machine and phased some cavities (not sure which ones).

Tried nudging beam around in the hall while watching the halo rates but didn't have any success, Fig 3. Decided to take beam away and check the main machine. Did some minor steering from basically 4R->AS, 8T, and 2C.

Fig 1: Comparison of rates before steering to after. Upstream is quiet, around 0-3, at 75 nA. Previously was up to 100 at 5 nA.

Fig 2: Final orbit after steering.

Downtime Incident CLOSED: Elevated Upstream Halo Rates

Log Number: 4258546Logbook(s): Body: IDDescriptionPeriodCause12758Title:
Elevated Upstream Halo Rates

Summary:
Upstream halo rate was elevated to the point it was interfering with the hall's BTA, similar to earlier in the week. Stopped to do some corrections throughout the main machine.

Downtime Incident CREATED: Elevated Upstream Halo Rates

Log Number: 4258545Logbook(s): Body: IDDescriptionPeriodCause12758Title:
Elevated Upstream Halo Rates

Summary:
Upstream halo rate was elevated to the point it was interfering with the hall's BTA, similar to earlier in the week. Stopped to do some corrections throughout the main machine.

FSD Masking Tool: Successful System UN-Masking for: BeamLossMonitor

Log Number: 4258539Logbook(s): Tags: Body: System UN-Masking for: BeamLossMonitor
User: deir
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T0

2C Steering Revisit

Log Number: 4258538Logbook(s): Entry Makers: deirBody: 

After nudging xB beam down on the lambertson

8T Steering

Log Number: 4258537Logbook(s): Entry Makers: deirBody: 

B&C on lambertson, nudged B down a bit

4th Pass Separation

Log Number: 4258532Logbook(s): Entry Makers: deirBody: 

Fig 1: As found

Linked ELOG is a check from 2/21, looks identical. However, nudged it a bit more toward the extracted side.

Pages